# Optimizing Attention with Mirror Descent: Generalized Max-Margin Token Selection

### v3

@@ 1 Introduction
Attention mechanisms (Bahdanau et al., 2014) have transformed natural language processing (NLP) and large language models (LLMs). Initially developed for encoder-decoder recurrent neural networks (RNNs), attention enables the decoder to focus on relevant input segments rather than relying solely on a fixed-length hidden state. This approach became fundamental in transformers (Vaswani et al., 2017), where attention layers—computing softmax similarities among input tokens—are the architecture’s backbone. Transformers have driven rapid advancements in NLP with models like BERT (Devlin et al., 2019) and ChatGPT (OpenAI, 2023), and have become the preferred architecture for generative modeling (Chen et al., 2021b; Ramesh et al., 2021), computer vision (Dosovitskiy et al., 2021; Radford et al., 2021), and reinforcement learning (Driess et al., 2023; Chen et al., 2021a). This has led to increased exploration of the mathematical foundations of attention’s optimization.
To understand optimization dynamics of attention mechanisms, Tarzanagh et al. (2024, 2023) studied the implicit bias of gradient descent (GD) in a binary classification setting with a fixed linear decoder. This bias refers to the tendency of GD to learn specific weight characteristics when multiple valid solutions exist. For example, in linear logistic regression on separable data, GD favors solutions aligned with the max-margin class separator (Soudry et al., 2018; Ji and Telgarsky, 2018). Similarly, Tarzanagh et al. (2023, 2024) propose a model resembling a hard-margin Support Vector Machine (SVM)—specifically, ($\ell_{p}$-AttSVM) with $p=2$—which maximizes the margin between optimal and non-optimal input tokens based on their softmax logits. These studies show that as training progresses, the combined key-query weights $W(k)$ of the model increasingly align with the locally optimal solution $W^{\alpha}_{\mathrm{mm}}$—the minimizer of ($\ell_{p}$-AttSVM) with $p=2$, where $k$ denotes the training iteration. Expanding on these insights, Vasudeva et al. (2024a) explores global directional convergence and the convergence rate of GD under specific conditions. Sheen et al. (2024) further extends these findings by relaxing assumptions about the convergence of regularized paths for the $(W_{K},W_{Q})$ parameterization of the key-query matrix, showing that gradient flow minimizes the nuclear norm of the key-query weight $W=W_{K}W_{Q}^{\top}$.
@@ 1.1 Contributions
While the aforementioned works provide insights into the implicit bias and token-selection properties of attention mechanisms, their analyses are limited to simplistic GD models. A broader understanding of more general descent algorithms, such as the mirror descent (MD) family, and their token-selection properties is missing. We address this by examining a family of MD algorithms designed for softmax attention, where the potential function is the $p$-th power of the $\ell_{p}$-norm, termed $\ell_{p}$-AttGD. This generalizes both $\ell_{p}$-GD (Azizan and Hassibi, 2018; Sun et al., 2022, 2023) and attention GD (Tarzanagh et al., 2024, 2023), enabling the exploration of key aspects of attention optimization via $\ell_{p}$-AttGD.
@@
Implicit bias of $\ell_{p}$-AttGD for attention optimization. Building on the setting of Tarzanagh et al. (2023); Vasudeva et al. (2024a); Sheen et al. (2024), we consider a one-layer attention model for binary classification. Specifically, given a dataset $(X_{i},y_{i},z_{i})_{i=1}^{n}$ where $X_{i}\in\mathbb{R}^{T\times d}$ represents inputs with $T$ tokens, $y_{i}\in\{\pm 1\}$ is the label, and $z_{i}\in\mathbb{R}^{d}$ is the comparison token, we study a single-layer attention model $f(X_{i},z_{i}):=v^{\top}X^{\top}_{i}\sigma(X_{i}Wz_{i})$, where $\sigma(\cdot)$ is the softmax function, $W$ is the key-query matrix, and $v$ is a linear decoder. Our goal is to separate a locally optimal token $\alpha_{i}\in[T]$ of each input sequence $X_{i}$ from the rest via an empirical risk minimization problem (ERM) with a smooth decreasing loss. We extend the SVM formulation of Tarzanagh et al. (2023) to ($\ell_{p}$-AttSVM), defining a hard-margin SVM using the $\ell_{p}$-norm instead of the $\ell_{2}$-norm. The solution $W^{\alpha}_{\mathrm{mm}}$ separates the locally optimal tokens $(\alpha_{i})_{i=1}^{n}$ with generalized maximum margin (see Section 2). Theorem 10 provides sufficient conditions for $\ell_{p}$-AttGD to converge locally, in direction, to $W_{\mathrm{mm}}^{\alpha}$. Moreover, Theorem 8 shows that $\|W(k)\|_{p,p}$ diverges as $k\rightarrow\infty$. These results characterize the implicit bias towards ($\ell_{p}$-AttSVM) in separating locally optimal tokens, extending previous work to a broader class of algorithms. While Theorem 10 and the results of Tarzanagh et al. (2024, 2023) offer insights into optimization dynamics for $p=2$, the finite-time convergence rate of $\ell_{p}$-AttGD for selecting locally optimal tokens remains unexplored.
@@
Convergence rate of $\ell_{p}$-AttGD to the solution of ($\ell_{p}$-AttSVM). Theorem 11 establishes convergence rates for $\ell_{p}$-AttGD, showing that the iterates $W(k)$, for large $k$, exhibit a decrease in $D_{\psi}\left(W_{\mathrm{mm}}^{\alpha}/\|W_{\mathrm{mm}}^{\alpha}\|_{p,p},W(k)/\|W(k)\|_{p,p}\right)$ at an inverse poly-log rate, where $D_{\psi}(\cdot,\cdot)$ denotes the Bregman divergence (Bregman, 1967b); see Definition 1. Despite optimizing a highly nonlinear, nonconvex softmax function, we achieve a convergence rate similar to GD in linear binary classification (Ji and Telgarsky, 2018, Theorem 1.1). Compared to the recent polynomial rate $O(k^{-3/4})$ in (Vasudeva et al., 2024a, Theorem 1) for optimizing attention, our rate is logarithmic and slower, but applicable to standard GD and MD for locally optimal token selection. Importantly, we do not require the near-orthogonality of tokens assumption used in Vasudeva et al. (2024a).
@@
Generalized Max-Margin Solutions and Joint Optimization of $(v,W)$. We investigate the joint problem over $v$ and $W$ under logistic loss using $\ell_{p}$-norm regularization path, where (ERM) is solved under $\ell_{p}$-norm constraints, examining the solutions trajectory as these constraints relax. Since the problem is linear in $v$, if the attention features $\bar{X}_{i}=X_{i}^{\top}\sigma(X_{i}Wz_{i})$ are separable by their labels $y_{i}$, $v$ acts as a generalized max-margin classifier (Azizan et al., 2021). Inspired by Tarzanagh et al. (2024, 2023), we show that under suitable geometric conditions, $W$ and $v$ generated by $\ell_{p}$-norm regularization path converge to their respective max-margin solutions (Theorem 31 in the appendix).
@@ 1.2 Additional Related Work
Oymak et al. (2023) and Deora et al. (2023) explore optimization dynamics in prompt-attention and multi-head attention models, respectively. Tian et al. (2023a, b) study SGD dynamics and multi-layer Transformer training. Tarzanagh et al. (2024, 2023) examine the implicit bias of GD in attention optimization. Vasudeva et al. (2024a) analyze the global directional convergence and convergence rate of GD for attention optimization under specific data conditions. Furthermore, Tarzanagh et al. (2023) and Sheen et al. (2024) respectively show that the regularization path and gradient flow not only achieve minimal loss but also minimize the nuclear norm of the key-query weight $W$. Thrampoulidis (2024); Li et al. (2024); Zhao et al. (2024) also study the optimization dynamics of attention mechanisms and the implicit bias of GD in next-token prediction. Our work extends these findings and those of Tarzanagh et al. (2024, 2023), focusing on the implicit bias of the general class of MD algorithms for attention training.
Implicit Bias of First-Order Methods. Significant progress has been made in understanding the implicit bias of gradient descent on separable data, particularly highlighted by Soudry et al. (2018); Ji and Telgarsky (2018). For linear predictors, Nacson et al. (2019); Ji and Telgarsky (2021); Ji et al. (2021) demonstrate that gradient descent methods rapidly converge to the max-margin predictor. Extending these insights to MLPs, Ji and Telgarsky (2020); Lyu and Li (2019); Chizat and Bach (2020) examine the implicit bias of GD and gradient flow using exponentially-tailed classification losses, showing convergence to the Karush-Kuhn-Tucker (KKT) points of the corresponding max-margin problem, both in finite (Ji and Telgarsky, 2020; Lyu and Li, 2019) and infinite-width scenarios (Chizat and Bach, 2020). Furthermore, the implicit bias of GD for training ReLU and Leaky-ReLU networks has been investigated, particularly in the context of orthogonal data (Phuong and Lampert, 2020; Frei et al., 2022). Additionally, the implicit bias toward rank minimization in regression settings with square loss has been explored in (Vardi and Shamir, 2021; Arora et al., 2019; Li et al., 2020).
@@ 2 Preliminaries
Let $N\geq 1$ and $[N]=\{1,2,\dots,N\}$. Vectors are denoted by lowercase letters (e.g., $a$), with components $a_{i}$, and matrices by uppercase letters (e.g., $A$). The minimum and maximum of scalars $a$ and $b$ are $a\wedge b$ and $a\vee b$, respectively. For a vector $v\in\mathbb{R}^{d}$, the $p$-norm is $\|v\|_{p}=(\sum_{i=1}^{d}|v_{i}|^{p})^{1/p}$. For a matrix $M\in\mathbb{R}^{d\times d}$, the $p,p$-norm is $\|M\|_{p,p}=(\sum_{i=1}^{d}\sum_{j=1}^{d}|M_{ij}|^{p})^{1/p}$. When $p=2$, these are the Euclidean norm for vectors and the Frobenius norm for matrices. For any two matrices $X,Y$ of the same dimensions, we define $\langle X,Y\rangle:=\text{trace}(X^{\top}Y)$. $\mathcal{O}$ denotes an upper bound, $\Omega$ a lower bound, and $\Theta$ both. All logarithms are natural (base $e$). Throughout, for a differentiable function $f:\mathbb{R}^{d\times d}\rightarrow\mathbb{R}$, we define $D_{f}:\mathbb{R}^{d\times d}\times\mathbb{R}^{d\times d}\rightarrow\mathbb{R}$ as
@@
where $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{1}}$, $W_{V}\in\mathbb{R}^{d\times d_{2}}$ are trainable key, query, value matrices, respectively; $\textnormal{softmax}(XW_{Q}W_{K}^{\top}Z^{\top})$ is the attention map; and $\textnormal{softmax}(\cdot):\mathbb{R}^{T\times T}\rightarrow\mathbb{R}^{T\times T}$ denotes the row-wise softmax function applied row-wise on $XW_{Q}W_{K}^{\top}Z^{\top}$. Similar to Tarzanagh et al. (2024, 2023), we reparameterize the key-query product matrix as $W:=W_{Q}W_{K}^{\top}\in\mathbb{R}^{d\times d}$, and subsume the value weights $W_{V}$ within the prediction head $v\in\mathbb{R}^{d}$. Suppose the first token of $Z$, denoted by $z$, is used for prediction. Then, the attention model can be formulated as
@@
Assumption A aligns with the assumptions on loss functions in Tarzanagh et al. (2024, 2023). Commonly used loss functions, such as $l(x)=\exp(-x)$, $l(x)=-x$, and $l(x)=\log(1+\exp(-x))$, satisfy this assumption.
Preliminaries on mirror descent. We review the MD algorithm (Blair, 1985) for solving attention-based (ERM). Mirror descent is defined using a potential function. We focus on differentiable and strictly convex potentials $\psi$ defined on the entire domain $\mathbb{R}^{d\times d}$. Note that in general, the potential function is a convex function of Legendre type (Rockafellar, 2015, Section 26). We call $\nabla\psi$ the mirror map. The natural “distance” associated with the potential $\psi$ is given by the Bregman divergence (Bregman, 1967a).
@@
Equivalently, MD can be written as $\nabla\psi(W(k+1))=\nabla\psi(W(k))-\eta\nabla\mathcal{L}(W(k))$; see Bubeck et al. (2015); Juditsky and Nemirovski (2011). A useful fact about the Bregman divergence is that it is non-negative and $D_{\psi}(W,V)=0$ if and only if $W=V$.
@@
$\begin{split}W^{\alpha}_{\mathrm{mm}}&:=\operatorname*{arg\,min}_{W\in\mathbb{R}^{d\times d}}\|W\|_{F}\\ &\textnormal{subj.\leavevmode\nobreak\ to}\leavevmode\nobreak\ \leavevmode\nobreak\ (X_{i\alpha_{i}}-X_{it})^{\top}Wz_{i}\geq 1,\leavevmode\nobreak\ \leavevmode\nobreak\ \textnormal{for all }\leavevmode\nobreak\ t\in[T]-\{\alpha_{i}\},\,i\in[n].\end{split}$ (3)
@@
It is worth noting that token scoring and optimal token identification can help us understand the importance of individual tokens and their impact on the overall objective. A token score measures how much a token contributes to a prediction or classification task, while an optimal token is defined as the token with the highest relevance in the corresponding input sequence (Tarzanagh et al., 2024, 2023). For illustration, please refer to Figure 1.

### v5

@@ 1 Introduction
Attention mechanisms (Bahdanau et al., 2014) have transformed natural language processing (NLP) and large language models (LLMs). Initially developed for encoder-decoder recurrent neural networks (RNNs), attention enables the decoder to focus on relevant input segments rather than relying solely on a fixed-length hidden state. This approach became fundamental in transformers (Vaswani et al., 2017), where attention layers—computing softmax similarities among input tokens—are the architecture’s backbone. Transformers have driven rapid advancements in NLP with models like BERT (Devlin et al., 2019) and ChatGPT (OpenAI, 2023), and have become the preferred architecture for generative modeling (Chen et al., 2021b; Ramesh et al., 2021), computer vision (Dosovitskiy et al., 2021b; Radford et al., 2021), and reinforcement learning (Driess et al., 2023; Chen et al., 2021a). This has led to increased exploration of the mathematical foundations of attention’s optimization.
To understand optimization dynamics of attention mechanisms, Tarzanagh et al. (2023, 2024) studied the implicit bias of gradient descent (GD) in a binary classification setting with a fixed linear decoder. This bias refers to the tendency of GD to learn specific weight characteristics when multiple valid solutions exist. For example, in linear logistic regression on separable data, GD favors solutions aligned with the max-margin class separator (Soudry et al., 2018; Ji and Telgarsky, 2018). Similarly, Tarzanagh et al. (2023, 2024) propose a model resembling a hard-margin Support Vector Machine (SVM)—specifically, ($\ell_{p}$-AttSVM) with p=2p=2—which maximizes the margin between optimal and non-optimal input tokens based on their softmax logits. These studies show that as training progresses, the combined key-query weights $W(k)$ of the model increasingly align with the locally optimal solution $W^{\alpha}_{\mathrm{mm}}$—the minimizer of ($\ell_{p}$-AttSVM) with $p=2$, where $k$ denotes the training iteration. Expanding on these insights, Vasudeva et al. (2024a) explores global directional convergence and the convergence rate of GD under specific conditions. Sheen et al. (2024) further extends these findings by relaxing assumptions about the convergence of regularized paths for the $(W_{K},W_{Q})$ parameterization of the key-query matrix, showing that gradient flow minimizes the nuclear norm of the key-query weight $W=W_{K}W_{Q}^{\top}$.
@@ 1.1 Contributions
While the aforementioned works provide insights into the implicit bias and token-selection properties of attention mechanisms, their analyses are limited to simplistic GD models. A broader understanding of more general descent algorithms, such as the mirror descent (MD) family, and their token-selection properties is missing. We address this by examining a family of MD algorithms designed for softmax attention, where the potential function is the $p$-th power of the $\ell_{p}$-norm, termed $\ell_{p}$-AttGD. This generalizes both $\ell_{p}$-GD (Azizan and Hassibi, 2018; Sun et al., 2022, 2023) and attention GD (Tarzanagh et al., 2023, 2024), enabling the exploration of key aspects of attention optimization via $\ell_{p}$-AttGD.
@@
Implicit Bias of $\ell_{p}$-AttGD for Attention Optimization. Building on the setting of Tarzanagh et al. (2023); Vasudeva et al. (2024a); Sheen et al. (2024), we consider a one-layer attention model for binary classification. Specifically, given a dataset $(X_{i},y_{i},z_{i})_{i=1}^{n}$ where $X_{i}\in\mathbb{R}^{T\times d}$ represents inputs with $T$ tokens, $y_{i}\in\{\pm 1\}$ is the label, and $z_{i}\in\mathbb{R}^{d}$ is the comparison token, we study a single-layer attention model $f(X_{i},z_{i}):=v^{\top}X^{\top}_{i}\sigma(X_{i}Wz_{i})$, where $\sigma(\cdot)$ is the softmax function, $W$ is the key-query matrix, and $v$ is a linear decoder. This single-head, one-layer setting is adopted for analytical tractability—a standard abstraction in theoretical studies of attention mechanisms (Tarzanagh et al., 2023, 2024; Vasudeva et al., 2024a; Sheen et al., 2024)—while our experimental validation in Section 4.2 demonstrates that these principles extend to practical multi-head, multi-layer architectures on real datasets. Our goal is to separate a locally optimal token $\alpha_{i}\in[T]$ of each input sequence $X_{i}$ from the rest via an empirical risk minimization problem (ERM) with a smooth decreasing loss. We extend the SVM formulation of Tarzanagh et al. (2023) to ($\ell_{p}$-AttSVM), defining a hard-margin SVM using the $\ell_{p}$-norm instead of the $\ell_{2}$-norm. The solution $W^{\alpha}_{\mathrm{mm}}$ separates the locally optimal tokens $(\alpha_{i})_{i=1}^{n}$ with generalized maximum margin (see Section 2).
Theorem 10 provides sufficient conditions for $\ell_{p}$-AttGD to converge locally, in direction, to $W_{\mathrm{mm}}^{\alpha}$. Moreover, Theorem 8 shows that $\|W(k)\|_{p,p}$ diverges as $k\rightarrow\infty$. These results characterize the implicit bias towards ($\ell_{p}$-AttSVM) in separating locally optimal tokens, extending previous work to a broader class of algorithms. While Theorem 10 and the results of Tarzanagh et al. (2023, 2024) offer insights into optimization dynamics for $p=2$, the finite-time convergence rate of $\ell_{p}$-AttGD for selecting locally optimal tokens remains unexplored.
@@
Convergence Rate of $\ell_{p}$-AttGD to the Solution of ($\ell_{p}$-AttSVM). Theorem 11 establishes convergence rates for $\ell_{p}$-AttGD, showing that the iterates $W(k)$, for large $k$, exhibit a decrease in $D_{\psi}\left(W_{\mathrm{mm}}^{\alpha}/\|W_{\mathrm{mm}}^{\alpha}\|_{p,p},W(k)/\|W(k)\|_{p,p}\right)$ at an inverse poly-log rate, where $D_{\psi}(\cdot,\cdot)$ denotes the Bregman divergence (Bregman, 1967); see Definition 1. Despite optimizing a highly nonlinear, nonconvex softmax function, we achieve a convergence rate similar to GD in linear binary classification (Ji and Telgarsky, 2018, Theorem 1.1). Compared to the recent polynomial rate $O(k^{-3/4})$ in (Vasudeva et al., 2024a, Theorem 1) for optimizing attention, our rate is logarithmic and slower, but applicable to standard GD and MD for locally optimal token selection. Importantly, we do not require the near-orthogonality of tokens assumption used in Vasudeva et al. (2024a).
@@
Generalized Max-Margin Solutions and Joint Optimization of $(v,W)$. We investigate the joint problem over $v$ and $W$ under logistic loss using $\ell_{p}$-norm regularization path, where (ERM) is solved under $\ell_{p}$-norm constraints, examining the solutions trajectory as these constraints relax. Since the problem is linear in $v$, if the attention features $\bar{X}_{i}=X_{i}^{\top}\sigma(X_{i}Wz_{i})$ are separable by their labels $y_{i}$, $v$ acts as a generalized max-margin classifier (Azizan et al., 2021). Inspired by Tarzanagh et al. (2023, 2024), we show that under suitable geometric conditions, $W$ and $v$ generated by $\ell_{p}$-norm regularization path converge to their respective max-margin solutions (Theorem 31 in the appendix).
@@ 1.2 Additional Related Work
Oymak et al. (2023) and Deora et al. (2023) explore optimization dynamics in prompt-tuning and multi-head attention models, respectively. Tian et al. (2023a, b) study SGD dynamics and multi-layer Transformer training. Tarzanagh et al. (2024, 2023) examine the implicit bias of GD in attention optimization. Vasudeva et al. (2024a) analyze the global directional convergence and convergence rate of GD for attention optimization under specific data conditions. Furthermore, Tarzanagh et al. (2023) and Sheen et al. (2024) respectively show that the regularization path and gradient flow not only achieve minimal loss but also minimize the nuclear norm of the key-query weight $W$. Thrampoulidis (2024); Li et al. (2024); Zhao et al. (2024) also study the optimization dynamics of attention mechanisms and the implicit bias of GD in next-token prediction. Our work extends these findings and those of Tarzanagh et al. (2023, 2024), focusing on the implicit bias of the general class of MD algorithms for attention training. Similar to these prior works, our theoretical analysis adopts a single-head, one-layer setting—a standard abstraction that balances analytical tractability with capturing the essential nonconvex optimization structure of softmax attention. Extensions to multi-head and multi-layer attention theory, as explored in (Tian et al., 2023b; Collins et al., 2024; Chen and Li, 2025), remain an important direction for future work.
Implicit Bias of First-Order Methods. Significant progress has been made in understanding the implicit bias of gradient descent on separable data, particularly highlighted by Soudry et al. (2018); Ji and Telgarsky (2018). For linear predictors, Nacson et al. (2019); Ji and Telgarsky (2021); Ji et al. (2021) demonstrate that gradient descent methods rapidly converge to the max-margin predictor. Extending these insights to multi-layer perceptrons (MLPs), Ji and Telgarsky (2020); Lyu and Li (2019); Chizat and Bach (2020) examine the implicit bias of GD and gradient flow using exponentially-tailed classification losses, showing convergence to the Karush-Kuhn-Tucker (KKT) points of the corresponding max-margin problem, both in finite (Ji and Telgarsky, 2020; Lyu and Li, 2019) and infinite-width scenarios (Chizat and Bach, 2020). Furthermore, the implicit bias of GD for training ReLU and Leaky-ReLU networks has been investigated, particularly in the context of orthogonal data (Phuong and Lampert, 2020; Frei et al., 2022). Additionally, the implicit bias toward rank minimization in regression settings with square loss has been explored in (Vardi and Shamir, 2021; Arora et al., 2019; Li et al., 2020).
@@ 2 Preliminaries
Let $N\geq 1$ and $N]=\{1,2,\dots,N\}$. Vectors are denoted by lowercase letters (e.g., $a$), with components $a_{i}$, and matrices by uppercase letters (e.g., $A$). The minimum and maximum of scalars $a$ and $b$ are $a\wedge b$ and $a\vee b$, respectively. For a vector $v\in\mathbb{R}^{d}$, the $p$-norm is $\|v\|_{p}=(\sum_{i=1}^{d}|v_{i}|^{p})^{1/p}$. For a matrix $M\in\mathbb{R}^{d\times d}$, the $p,p$-norm is $\|M\|_{p,p}=(\sum_{i=1}^{d}\sum_{j=1}^{d}|M_{ij}|^{p})^{1/p}$. When $p=2$, these are the Euclidean norm for vectors and the Frobenius norm for matrices. For any two matrices $X,Y$ of the same dimensions, we define $\langle X,Y\rangle:=\text{trace}(X^{\top}Y)$. $\mathcal{O}$ denotes an upper bound, $\Omega$ a lower bound, and $\Theta$ both. All logarithms are natural (base $e$). Throughout, for a differentiable function $f:\mathbb{R}^{d\times d}\rightarrow\mathbb{R}$, we define $D_{f}:\mathbb{R}^{d\times d}\times\mathbb{R}^{d\times d}\rightarrow\mathbb{R}$ as
@@
where $W_{Q},W_{K}\in\mathbb{R}^{d\times d_{1}}$, $W_{V}\in\mathbb{R}^{d\times d_{2}}$ are trainable key, query, value matrices, respectively; $\textnormal{softmax}(XW_{Q}W_{K}^{\top}Z^{\top})$ is the attention map; and $\textnormal{softmax}(\cdot):\mathbb{R}^{T\times T}\rightarrow\mathbb{R}^{T\times T}$ denotes the row-wise softmax function applied row-wise on $XW_{Q}W_{K}^{\top}Z^{\top}$. Similar to Tarzanagh et al. (2023, 2024), we reparameterize the key-query product matrix as $W:=W_{Q}W_{K}^{\top}\in\mathbb{R}^{d\times d}$, and subsume the value weights $W_{V}$ within the prediction head $v\in\mathbb{R}^{d}$. Suppose the first token of $Z$, denoted by $z$, is used for prediction. Then, the attention model can be formulated as
@@
Assumption A aligns with the assumptions on loss functions in Tarzanagh et al. (2024, 2023). Commonly used loss functions, such as $l(x)=\exp(-x)$, l(x)=xl(x)=-x, and $l(x)=\log(1+\exp(-x))$, satisfy this assumption.
Preliminaries on mirror descent. We review the MD algorithm (Blair, 1985) for solving attention-based (ERM). Mirror descent is defined using a potential function. We focus on differentiable and strictly convex potentials $\psi$ defined on the entire domain $\mathbb{R}^{d\times d}$. Note that in general, the potential function is a convex function of Legendre type (Rockafellar, 2015, Section 26). We call $\nabla\psi$ the mirror map. The natural “distance” associated with the potential $\psi$ is given by the Bregman divergence (Bregman, 1967).
@@
Equivalently, MD can be written as $\nabla\psi(W(k+1))=\nabla\psi(W(k))-\eta\nabla\mathcal{L}(W(k))$; see Bubeck and others (2015); Juditsky and Nemirovski (2011). A useful fact about the Bregman divergence is that it is non-negative and $D_{\psi}(W,V)=0$ if and only if $W=V$.
@@
$\begin{split}W^{\alpha}_{\mathrm{mm}}&:=\operatorname*{arg\,min}_{W\in\mathbb{R}^{d\times d}}\|W\|_{F}\\ &\textnormal{subj.\penalty 10000\ to}\penalty 10000\ \penalty 10000\ (X_{i\alpha_{i}}-X_{it})^{\top}Wz_{i}\geq 1,\penalty 10000\ \penalty 10000\ \textnormal{for all }\penalty 10000\ t\in[T]-\{\alpha_{i}\},\,i\in[n].\end{split}$ (3)
@@
It is worth noting that token scoring and optimal token identification can help us understand the importance of individual tokens and their impact on the overall objective. A token score measures how much a token contributes to a prediction or classification task, while an optimal token is defined as the token with the highest relevance in the corresponding input sequence (Tarzanagh et al., 2023). For illustration, please refer to Figure 1.